OpenStudy (ammarah):

Solve the equation. Round the result to two decimal places when appropriate. x^2=59

OpenStudy (compassionate):

Square x on both sides.

OpenStudy (cwrw238):

take the square root of each side

OpenStudy (ammarah):

oh wait i messed up the equation. its x^-5=59

OpenStudy (ammarah):

so it would be x=8? or -8

OpenStudy (ammarah):

@Compassionate please see if this is right answer

OpenStudy (cwrw238):



OpenStudy (ammarah):

noo its x^2-5

OpenStudy (ammarah):

omg im trying to look for someone to check it...

OpenStudy (anonymous):

x^2-5=59 add 5 both sides x^2=64 take square root both sides x=sqrt(64) x=8,-8 because 8^2=64 and -8^2=64

OpenStudy (ammarah):

I need help with this oene too: